logarithms confuse me..
solve: ln(x)-ln(x-4)=ln(3)

Answer by edjones(8007)

ln(x)-ln(x-4)=ln(3)
e^(ln(x))/e^(ln(x-4))=e^(ln(3))
x/(x-4)=3
x=3x-12
2x=12
x=6
.
Ed